OpenStudy (mayah3):

Choose the correct description of the graph of the compound inequality x - 3 < -7 or x + 5 greater than or equal to 8

OpenStudy (mayah3):

Open circle on 8, shading to the left. Closed circle on 8, shading to the left. Open circle on 8, shading to the right. Closed circle on 8, shading to the right.

OpenStudy (shamil98):

x - 3 < -7 , x + 5 ≥ 8 solve for x and write it as a compound inequality. then graph it on the number line.

OpenStudy (mayah3):

.. you just restated the question..

OpenStudy (shamil98):

I just told you how to do it. This site isn't here to do your work.

OpenStudy (shamil98):

x - 3 < -7 , x + 5 ≥ 8 is not written in a compound inequality form the compound inequality form would be something like 3 < x ≤ 8

OpenStudy (shamil98):

From there it should be fairly simple to graph it on the number line.

OpenStudy (mayah3):

If I knew how to dove for x I would have done it. I'm not asking for someone to do it. I'm asking for understanding of the topic.

OpenStudy (anonymous):

what are the options?

OpenStudy (mayah3):

solve*

OpenStudy (mayah3):

Open circle on 8, shading to the left. Closed circle on 8, shading to the left. Open circle on 8, shading to the right. Closed circle on 8, shading to the right

OpenStudy (anonymous):

ok hold on and I'll try to explain

OpenStudy (mayah3):

Ok tyt

OpenStudy (anonymous):

first split it up x-3<-7 and x+5>=8 x-3<-7 +3 +3 x<-4 for the second part x+5>=8 -5 -5 x>=3 so it would end up being x is less than -4 and greater than or equal to 3 what do you think it is from there?

OpenStudy (mayah3):

You'd have to graph it?

OpenStudy (anonymous):

yes and I did something wrong hold on

OpenStudy (anonymous):

Open=Greater or less then Clossed= greater and equal to and less then and equal to.

OpenStudy (mayah3):

Oh okay holding on lol

OpenStudy (anonymous):

from that information what do you think it would be?

OpenStudy (mayah3):

would it be closed circle at -4 shading to the left and open circle at 3 shading to this right?

OpenStudy (anonymous):

I did something wrong on that so forget what I said on that lol from 8 would it be open or closed to start with?

OpenStudy (anonymous):

ok now I'm confused on the 8 thing hold on and I'm going to look it up

OpenStudy (mayah3):

Alright

OpenStudy (anonymous):

everything I have seen you simplify and go from there but this goes from 8 not the simplified x :/ we are supposed to be graphing x right?

OpenStudy (mayah3):

I honestly have no idea :/

OpenStudy (anonymous):

ok I see what they are doing. It will be closed since it is greater than or equal to x so it is either B or D. do you think the shading goes to the left or right?

OpenStudy (mayah3):

To the right.

OpenStudy (anonymous):

yes so D and sorry for the confusion

OpenStudy (anonymous):

I read it wrong and it made it really hard to understand but it should be D tell me if that's right (:

OpenStudy (mayah3):

Thank you, I will inform you if it is right (: